OpenStudy (anonymous):

The rectangles shown are similar. What is the value of x? A. 6 B. 4 C. 9 D. 5

OpenStudy (anonymous):

do you have a picture of the rectangles?

OpenStudy (anonymous):

it wont let me put them up

OpenStudy (anonymous):

ill try to draw it

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

here

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

hey can you help me with a few more if not its cool

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

What is the ratio of x to y? A. B. C. D.

OpenStudy (anonymous):

there is nothing there

OpenStudy (anonymous):

sorry took me a minute computers dum

OpenStudy (anonymous):

a -7x=y

OpenStudy (anonymous):

b y/x=-7

OpenStudy (anonymous):

c x/y=7

OpenStudy (anonymous):

d xy=-7

OpenStudy (funnyguy2):

x=-2y/3 y=-3x/2

OpenStudy (funnyguy2):

can you figure out the rest?

OpenStudy (anonymous):

flutter you

OpenStudy (anonymous):

wingspan

OpenStudy (anonymous):

noobs

OpenStudy (rock_mit182):

I would say for the first one question the correct answer is D

OpenStudy (anonymous):

why you on this question?